66 2/3 percent of 2kg, 6g, 3dag, 7g
Answer by CubeyThePenguin(3113)

2/3 * 2 kg = 4/3 kg

2/3 * 6 g = 4 g

2/3 * 6 dag = 4 dag

2/3 * 7 g = 14/3 g